# MIGC++: Advanced Multi-Instance Generation Controller for Image Synthesis

### v1

@@ 2 Related Work
Text-to-image generation aims to create images from textual descriptions. Initially, conditional Generative Adversarial Networks (GANs) [46, 47, 48, 49] were used for this purpose. However, diffusion models [50, 23, 36, 51, 52, 53, 54, 38, 55, 56, 57, 58, 59] and autoregressive models [60, 61, 62] have largely replaced GANs, offering more stable training and enhanced image quality. To control generated content, Guided Diffusion [63] utilizes classifiers to assess generated images and guidances the sampling through the gradient of the classifiers. Classifier-free guidance (CFG) [54] further optimizes this process by interpolating between conditioned and unconditioned predictions, with methods like GLIDE [50], showing promising results under these protocols. DALL-E 2 [53] employs transformations from CLIP text features [24] to CLIP image space for image generation, while Imagen [36] leverages the T5 large language model [64] as a text encoder to improve quality. eDiff-I [52] represents a significant progression in the field by integrating an expert generator to elevate image quality. Stable Diffusion [23], as one open-source model, is one of the most widely studied Text-to-Image models.
@@ 4.4 Consistent-MIG
TABLE II: Quantitative results (§6.1) on proposed COCO-MIG-MASK.
@@ 6 Experimental Results
We assessed MIGC/MIGC++ using our COCO-MIG and Multimodal-MIG benchmarks and other established ones.:
@@ 6.1 Compare with State-of-the-art Competitors
COCO-MIG-MASK. As illustrated in Tab. II, compared to Instance Diffusion  [26], MIGC++ has improved by 16% in Instance Success Ratio, 12.5% in Image Success Ratio, and 13.3% in MIoU. Particularly, we found that as the number of instances increases, MIGC++ achieves more accurate Multi-Instance Generation than Instance Diffusion. For example, when the number of instances increases from 2 to 6, the improvement of MIGC++ over Instance Diffusion increases from 8.5% to 17.8%. Qualitative comparisons in Fig.12 further elucidate these metrics, where MIGC++ ensures each instance accurately follows the shapes specified by the masks and has the correct color attributes. As a plug-and-play controller, we also demonstrate the effectiveness of MIGC++ on RV6.0 (i.e., a model fine-tuned from SD1.5), which can generate more realistic images.
@@ 6.2 Ablation Study
Inhibition Loss. Tab. IX indicates that using an inhibition loss with a weight of 0.1, as compared to not using this loss, allows MIGC to increase its AP by 2.23 and MIGC++ by 2.08. Increasing the loss weight to 1.0 further improves AP but at the expense of FID performance, leading to the selection of a 0.1 weight. Fig. 18, reveals that the advantages of inhibition loss are more pronounced with higher numbers of generated instances. For instance, with two instances, inhibition loss only improves MIoU from 65.3 to 65.8 (a 0.5 increase), but with five instances, it boosts from 55.4 to 57.5 (a 2.1 increase). Comparing Fig. 17(d) and Fig. 17(f), it is apparent that inhibition loss more effectively restricts the instances to their intended locations.
TABLE IX: Ablation Study of the proposed inhibition loss on the COCO-Position benchmark.
@@
Figure 18: Ablation study of the Inhibition Loss on the COCO-MIG benchmark (§6.2).

### v2

@@ 2 Related Work
Text-to-image generation aims to create images from textual descriptions. Initially, conditional Generative Adversarial Networks (GANs) [46, 47, 48, 49] were used for this purpose. However, diffusion models [50, 23, 36, 51, 52, 53, 54, 38, 55, 56, 57, 58, 59] and autoregressive models [60, 61, 62] have largely replaced GANs, offering more stable training and enhanced image quality. To control generated content, Guided Diffusion [63] utilizes classifiers to assess generated images and guidances the sampling through the gradient of the classifiers. Classifier-free guidance (CFG) [54] further optimizes this process by interpolating between conditioned and unconditioned predictions, with methods like GLIDE [50], showing promising results under these protocols. DALL-E 2 [53] employs transformations from CLIP text features [24] to CLIP image space for image generation, while Imagen [36] leverages the T5 large language model [64] as a text encoder to improve quality. eDiff-I [52] further elevates image quality by integrating expert generators. Stable Diffusion [23] incorporates the attention mechanism to seamlessly integrate textual information into the generation process.
@@ 4.4 Consistent-MIG
TABLE II: Quantitative results (§6.1) on proposed COCO-MIG-MASK. : Using only box annotation data during training.
@@ 6 Experimental Results
We assessed MIGC/MIGC++ using our COCO-MIG and Multimodal-MIG benchmarks and other established ones:
@@ 6.1 Compare with State-of-the-art Competitors
COCO-MIG-MASK. As illustrated in Tab. II, MIGC++ extends the application of MIGC by allowing masks to describe the fine-grained positions of instances, achieving a 17.2% improvement in MIoU. Compared to Instance Diffusion  [26], MIGC++ has improved by 16% in Instance Success Ratio, 12.5% in Image Success Ratio, and 13.3% in MIoU. Qualitative comparisons in Fig.12 further elucidate these metrics, where MIGC++ ensures each instance accurately follows the shapes specified by the masks and has the correct color attributes. As a plug-and-play controller, we also demonstrate the effectiveness of MIGC++ on RV6.0 (i.e., a model fine-tuned from SD1.5), which can generate more realistic images.
@@
Robustness to Increasing Instance Number. Results from Tab. I and Tab. II reveal that MIGC++ maintains more consistent performance compared to previous methods as the number of instances increases, thanks to two key enhancements: 1) The divide-and-conquer framework allows MIGC++ to break down multi-instance generation into independent subtasks, which minimizes the impact on every single instance as the number of instances grows—e.g., while InstanceDiffusion saw a 16.5% ISR drop from L2 to L6, MIGC++ experienced only a 6.2% ISR drop in Tab. II. 2) The introduction of a refined shader, replacing the remaining Cross-Attention layers, secures independent shading of each instance throughout the network’s forward process—where MIGC’s performance decreased by 8.5% from L2 to L6, MIGC++ saw a drop of just 3.3% in Tab. I.
@@ 6.2 Ablation Study
Inhibition Loss. Tab. IX indicates that using an inhibition loss with a weight of 0.1, as compared to not using this loss, allows MIGC to increase its AP by 2.23 and MIGC++ by 2.08. Although further increasing the loss weight offers a slight boost in control ability, it somewhat impacts the quality of generated images. For instance, with a higher loss weight (i.e., 1.0), the FID score worsens by 2.42 for MIGC and by 0.74 for MIGC++. Since both control ability and image quality are crucial for generative models, we set the default inhibition loss weight to 0.1 to enhance control without affecting image quality. Fig. 18, reveals that the advantages of inhibition loss are more pronounced with higher numbers of generated instances. For instance, with two instances, inhibition loss only improves MIoU from 65.3 to 65.8 (a 0.5 increase), but with five instances, it boosts from 55.4 to 57.5 (a 2.1 increase). Comparing Fig. 17(d) and Fig. 17(f), it is apparent that inhibition loss more effectively restricts the instances to their intended locations.
[…]
@@
Figure 18: Ablation study of the Inhibition Loss on the COCO-MIG benchmark (§6.2). We investigate the improvements brought by inhibition loss across different instance counts using the default loss weight (see Tab. IX).